I was wondering if you still answer questions because i have not received and answer in a few days

these were the question. it says use the substitution method to solve the linear system

1. 3x-2y =4
x+3y =5

2. 3x+y = - 2
x + 3y = 2

THANK YOU IF YOU CAN HELP

Answer by AnlytcPhil(1806)

I was wondering if you still answer questions because
i have not received and answer in a few days

==========================

You'll probably recievs a faster reply if you post only
ONE problem at a time.

There are different ways of solving these. This method
is the substitution method.  If you want the addition
method, post again and state this.

1.      3x - 2y = 4
         x + 3y = 5 

Solve the second equation for x

         x + 3y = 5
              x = 5 - 3y
 
Substitute (5-3y) for x in the first equation
 
        3x - 2y = 4
   3(5-3y) - 2y = 4 
   15 - 9y - 2y = 4
       15 - 11y = 4
           -11y = 4 - 15
           -11y = -11
              y = -11/(-11)
              y = 1 

Substitute (1) for y in this equation where you
solved for x above:

              x = 5 - 3y  
              x = 5 - 3(1)
              x = 5 - 3
              x = 2 

2.      3x +  y = -2
         x + 3y =  2

Solve the first equation for y

         3x + y = -2
              y = -2 - 3x
 
Substitute (-2-3x) for y in the second equation
 
         x + 3y = 2
   x + 3(-2-3x) = 2 
     x - 6 - 9x = 2
        -6 - 8x = 2
            -8x = 2 + 6
            -8x = 8
              x = 8/(-8)
              x = -1 

Substitute (-1) for x in this equation where you
solved for y above:
              y = -2 - 3x
              y = -2 - 3(-1)  
              y = -2 - 3(-1)
              y = -2 + 3
              y = 1

Edwin McCravy
AnlytcPhil@aol.com